OpenStudy (anonymous):

Express answers in simplest exact form. What is the area of a sector with measure of arc equal to 90° and radius equal to 1 foot? A = ? PI sq. ft.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

\[\frac{90{}^{\circ}}{360{}^{\circ}}\left(\pi 1^2\right)=\frac{\pi }{4} \text{ ft}^2 \]

OpenStudy (anonymous):

lol man i dont get what You wrote lol i juss need the numbers lol